# CaLiV: LiDAR-to-Vehicle Calibration of Arbitrary Sensor Setups

### v2

@@ Abstract
In autonomous systems, sensor calibration is essential for safe and efficient navigation in dynamic environments. Accurate calibration is a prerequisite for reliable perception and planning tasks such as object detection and obstacle avoidance. Many existing LiDAR calibration methods require overlapping fields of view, while others use external sensing devices or postulate a feature-rich environment. In addition, Sensor-to-Vehicle calibration is not supported by the vast majority of calibration algorithms. In this work, we propose a novel target-based technique for extrinsic Sensor-to-Sensor and Sensor-to-Vehicle calibration of multi-LiDAR systems called CaLiV. This algorithm works for non-overlapping fields of view and does not require any external sensing devices. First, we apply motion to produce field of view overlaps and utilize a simple Unscented Kalman Filter to obtain vehicle poses. Then, we use the Gaussian mixture model-based registration framework GMMCalib to align the point clouds in a common calibration frame. Finally, we reduce the task of recovering the sensor extrinsics to a minimization problem. We show that both translational and rotational Sensor-to-Sensor errors can be solved accurately by our method. In addition, all Sensor-to-Vehicle rotation angles can also be calibrated with high accuracy. We validate the simulation results in real-world experiments. The code is open-source and available on https://github.com/TUMFTM/CaLiV.
© This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ III-B Rough Point Cloud Alignment
$\begin{split}{}_{L^{i}_{a}}^{C}T={}_{L^{i}_{a}}^{C}T^{{}^{\prime}}~{}_{V^{i}}^{R}T~{}_{L_{a}}^{V}\hat{T}\quad\quad\quad\\ \forall i,a\quad s.t.\quad 1\leq i\leq N,~a\in\{1,2\}.\end{split}$ (2)
$_{L^{i}_{a}}^{C}T^{{}^{\prime}}$ here is the part of the alignment that is done by the point cloud registration algorithm of GMMCalib and $a$ represents the specific LiDAR sensor.
@@ III-D Implementation Details
We now present the implementation details of our framework used to solve the minimization problem formulated in Section III-C. Specifically, we discuss the kinematic model used to recover the vehicle poses $_{V^{i}}^{R}T$, the joint registration algorithm used to obtain the transformations $_{L^{i}_{a}}^{C}T^{{}^{\prime}}$ of all point clouds to a common calibration frame $\{C\}$, as well as the minimization algorithm.
@@ III-D1 Kinematic Model
The approach is compatible with any kinematic model, provided that the vehicle poses are given at the time steps of the LiDAR measurements. We use an Unscented Kalman Filter (UKF) for the kinematic model that combines IMU, GPS, velocity and orientation data to recover vehicle pose estimates $_{V^{i}}^{R}T$ at all time steps $t_{i}$. Our UKF also allows the modeling of sensor-specific noise and uncertainties to evaluate the calibration accuracy under noisy sensor data.
@@ III-D2 Joint Registration
Before a point cloud registration algorithm can be applied to the point clouds, they must be pre-processed. First, the point clouds are transformed with the initial LiDAR transformations and the ego vehicle pose at the measurement’s time step, as discussed in Section III-B. Then, we separate the target points from the ground points by using the Random Sample Consensus (RANSAC) [26] algorithm to identify and filter out the ground plane. Finally, we use GMMCalib [10] for joint registration of the filtered point clouds into a common calibration frame $\{C\}$. GMMCalib returns the registration transform $_{L^{i}_{1}}^{C}T^{{}^{\prime}}$ and $_{L^{i}_{2}}^{C}T^{{}^{\prime}}$ for each point cloud that contains points belonging to the target shape.
@@ IV EXPERIMENT DESIGN
In this chapter, we present the experimental setup used to evaluate our proposed framework. The simulation is performed in the driving simulator CARLA [28] and consists of the digital twin of the real-world vehicle, a generic calibration target, and a flat surface. To validate our approach, the experiments are also performed in a real-world environment. Furthermore, we describe the evaluation metrics of the S2S and S2V calibration performance.
@@ IV-B Evaluation
The S2S calibration error can now be defined as the residual of the optimized $^{V}_{L_{1}}T^{*}$ (which still contains the S2S error) relative to the expression above, since both share the same S2V error. This way, the S2S error can be evaluated independently from the S2V error.
In our experiments, we added randomly generated errors to the ground truth transformations of $L_{1}$ and $L_{2}$ to obtain different initial errors on the calibration matrices. The translation errors for x, y and z are uniformly distributed in the range of 0.1m-0.1m to $0.1m$. The rotation errors for roll, pitch and yaw are also uniformly distributed, but in the range of $3.0\degree$ to $3.0\degree$. In total, we run 100 simulations containing noisy UKF poses.
@@ V-A S2S Calibration
The adaptive voxelization-based work of [23] produces rather inaccurate calibration transformations in our experiments. The mean translation errors are up to several meters, while the mean rotation errors are in the range of $6\degree$ to $19\degree$. While mlcc works very well when given accurate initial extrinsics and poses, it does not provide accurate results when exposed to erroneous LiDAR poses that are due to calibration errors. In their work, Liu et al. state that the initial LiDAR extrinsic parameters should be pre-calibrated and accurate LiDAR poses must available, e.g. by first applying a hand-eye calibration or SLAM. From our results, we conclude that the initial S2V errors of up to $3\degree$ introduced in our experiment cannot be handled by mlcc and a more accurate prior calibration is necessary.
@@ V-C Observability
We want to highlight the accuracy of the calibration with respect to the rotational components, as well as discuss the translational components. In our experiments, we found that the z-translation component is not observable with a planar trajectory. However, the z-translation can easily be compensated for using ground plane information, as presented in [16]. Another approach to handling the z-translation was proposed by iCalib [29], which used a physical ramp to introduce the necessary excitation in that dimension. More notably, due to the nature of the trajectory, the algorithm cannot distinguish between translational $x$-error and yaw error. In experiments, where the translation errors were not constrained to zero, we observed a calibration error for the reference sensor of $\delta x=0.2792~m$ and $\delta\phi=0.0304~rad$. With constrained translation errors the yaw angle can reliably be estimated with the accuracy reported in Table III. This could also be solved by driving longer trajectories, as demonstrated in [5]. However, the target-based approach limits the length and topology of the trajectory. This creates a trade-off in which the method prioritizes compensating for rotational calibration errors. It should be noted that the error of the $y$-component of the translation vector is not subject to the observability issues and achieves a mean accuracy of 0.0160m0.0160~m.
@@ V-D Real World
It should be noted that no prior sensor calibration was performed, resulting in relatively high initial calibration errors as shown on the left of Fig. 7. The figure shows the first frame of the front LiDAR at the beginning of the trajectory and the last frame of the rear LiDAR at the end of the trajectory. This gives a better insight into the overall ability of CaLiV to compensate for the propagating S2V calibration error over time/distance. After performing the calibration with CaLiV, the validation target seen on the right side of Fig. 7 is well aligned, indicating a successful S2V calibration.
To achieve these results, a number of criteria must be met in the real experiments. To account for the motion distortion of the LiDAR sensors and the distortion of the angular velocities during the curved trajectory, the vehicle motion must be stationary and at a constant velocity of about 5 $\frac{m}{s}$. Furthermore, real-world LiDAR sensors do not have homogeneous point resolution across the vertical and horizontal FoV, resulting in sparse point clouds at the edges of the FoV. In our experiments, this required an additional voxel and outlier filter to account for the differences in cardinality and improve the registration results during calibration.
V-E Computation Time
It is important to note that CaLiV cannot be used in real time, but rather as an offline algorithm. The current computation time of the algorithm is about 20 minutes, mainly due to the optimization step.
@@ VI CONCLUSION & FUTURE WORK
The goal of this work was to provide a target-based framework for S2S and S2V calibration of multi-LiDAR systems with arbitrary FoV overlap. In addition, we aimed to develop an algorithm that is independent of environmental features and does not rely on external sensors. CaLiV demonstrates high accuracy and robustness for both S2S and S2V calibration. CaLiV outperforms the current state of the art in both non-overlapping sensor calibration and S2V calibration. While our current method requires a curved trajectory to optimize S2S and S2V calibration, it remains to be investigated whether moving the target would yield equally accurate results. Furthermore, the possibility of using a more complex trajectory could be explored to also solve the S2V translation errors while maintaining the accuracy.

### v3

@@ Abstract
In autonomous systems, sensor calibration is essential for safe and efficient navigation in dynamic environments. Accurate calibration is a prerequisite for reliable perception and planning tasks such as object detection and obstacle avoidance. Many existing LiDAR calibration methods require overlapping fields of view, while others use external sensing devices or postulate a feature-rich environment. In addition, Sensor-to-Vehicle calibration is not supported by the vast majority of calibration algorithms. In this work, we propose a novel target-based technique for extrinsic Sensor-to-Sensor and Sensor-to-Vehicle calibration of multi-LiDAR systems called CaLiV. This algorithm works for non-overlapping fields of view and does not require any external sensing devices. First, we apply motion to produce field of view overlaps and utilize a simple Unscented Kalman Filter to obtain vehicle poses. Then, we use the Gaussian mixture model-based registration framework GMMCalib to align the point clouds in a common calibration frame. Finally, we reduce the task of recovering the sensor extrinsics to a minimization problem. We show that both translational and rotational Sensor-to-Sensor errors can be solved accurately by our method. In addition, all Sensor-to-Vehicle rotation angles can also be calibrated with high accuracy. We validate the simulation results in real-world experiments. The code is open source and available on https://github.com/TUMFTM/CaLiV.
@@ III-B Rough Point Cloud Alignment
$\begin{split}{}_{L^{i}_{a}}^{C}T={}_{R}^{C}\hat{T}~{}_{V^{i}}^{R}T~{}_{L_{a}}^{V}\hat{T}\quad\quad\quad\\ \forall i,a\quad s.t.\quad 1\leq i\leq N,~a\in\{1,2\}.\end{split}$ (2)
$_{R}^{C}\hat{T}$ here is the part of the alignment that is done by the point cloud registration algorithm of GMMCalib and $a$ represents the specific LiDAR sensor. Similar to the ICP algorithm, the registration of GMMCalib requires a good initial alignment of the point clouds as it positively affects the calibration results. This is achieved by the pre-alignment step discussed above.
@@ III-D Implementation Details
We now present the implementation details of our framework used to solve the minimization problem formulated in Section III-C. Specifically, we discuss the kinematic model used to recover the vehicle poses $_{V^{i}}^{R}T$, the joint registration algorithm used to obtain the transformations $_{L^{i}_{a}}^{C}T$ of all point clouds to a common calibration frame $\{C\}$, as well as the minimization algorithm.
@@ III-D1 Kinematic Model
The approach is compatible with any kinematic model, provided that the vehicle poses are given at the time steps of the LiDAR measurements. We use an Unscented Kalman Filter (UKF) for the kinematic model that combines IMU, GPS, velocity and orientation data to recover vehicle pose estimates $_{V^{i}}^{R}T$ at all time steps $t_{i}$. We assume these sensors to be calibrated. Our UKF also allows the modeling of sensor-specific noise and uncertainties to evaluate the calibration accuracy under noisy sensor data.
@@ III-D2 Joint Registration
Before a point cloud registration algorithm can be applied to the point clouds, they must be pre-processed. First, the point clouds are transformed with the initial LiDAR transformations and the ego vehicle pose at the measurement’s time step, as discussed in Section III-B. Then, we separate the target points from the ground points by using the Random Sample Consensus (RANSAC) [9] algorithm to identify and filter out the ground plane. Finally, we use GMMCalib [23] for joint registration of the filtered point clouds into a common calibration frame $\{C\}$. GMMCalib returns the registration transform $_{L^{i}_{1}}^{C}T$ and $_{L^{i}_{2}}^{C}T$ for each point cloud that contains points belonging to the target shape.
@@ IV EXPERIMENT DESIGN
In this chapter, we present the experimental setup used to evaluate our proposed framework. The simulation is performed in the driving simulator CARLA [8] and consists of the digital twin of the real-world vehicle EDGAR [13], a generic calibration target, and a flat surface. To validate our approach, the experiments are also performed in a real-world environment. Furthermore, we describe the evaluation metrics of the S2S and S2V calibration performance.
@@ IV-B Evaluation
The S2S calibration error can now be defined as the residual of the optimized $^{V}_{L_{1}}T^{*}$ (which still contains the S2S error) relative to the expression above, since both share the same S2V error. This way, the S2S error can be evaluated independently from the S2V error. In our experiments, we added randomly generated errors to the ground truth transformations of $L_{1}$ and $L_{2}$ to obtain different initial errors on the calibration matrices. The translation errors for x, y and z are uniformly distributed in the range of 0.1m-0.1m to $0.1m$. The rotation errors for roll, pitch and yaw are also uniformly distributed, but in the range of $3.0\degree$ to $3.0\degree$. In total, we run 100 simulations containing noisy UKF poses.
@@ V-A S2S Calibration
The mlcc framework [18] is an accurate method for non-overlapping S2S calibration. However, in our experimental setup, it yielded significantly less accurate results, with mean translation errors reaching several meters and mean rotation errors ranging from $6^{\circ}$ to $19^{\circ}$. While mlcc works very well when given accurate initial extrinsics and poses, it does not provide accurate results when exposed to erroneous LiDAR poses that are due to S2V calibration errors. In their work, Liu et al. state that the initial LiDAR extrinsic parameters should be pre-calibrated and accurate LiDAR poses must be available, e.g. by first applying a hand-eye calibration or SLAM. From our results, we conclude that the initial S2V errors of up to $3\degree$ introduced in our experiment cannot be handled by mlcc. Therefore, in contrast to CaLiV, a more accurate prior calibration is necessary.
@@ V-C Observability
We want to highlight the accuracy of the calibration with respect to the rotational components, as well as discuss the translational components. In our experiments, we found that the z-translation component is not observable with a planar trajectory. However, the z-translation can easily be compensated for using ground plane information, as presented in [14]. Another approach to handling the z-translation was proposed by iCalib [28], which used a physical ramp to introduce the necessary excitation in that dimension.
More notably, due to the nature of the trajectory, the algorithm cannot distinguish between translational $x$-error and yaw error. In experiments, where the translation errors were not constrained to zero, we observed a calibration error for the reference sensor of $\delta x=0.2792~m$ and $\delta\phi=0.0304~rad$. With constrained translation errors the yaw angle can reliably be estimated with the accuracy reported in Table III. This could also be solved by driving longer trajectories, as demonstrated in [4]. However, the target-based approach limits the length and topology of the trajectory. This creates a trade-off in which the method prioritizes compensating for rotational calibration errors. It should be noted that the error of the $y$-component of the translation vector is not subject to the observability issues and achieves a mean accuracy of 0.0160m0.0160~m.
@@ V-D Real World
It should be noted that no prior sensor calibration was performed, resulting in relatively high initial calibration errors as shown on the left of Fig. 7. The figure shows the first frame of the front LiDAR at the beginning of the trajectory and the last frame of the rear LiDAR at the end of the trajectory. This gives a better insight into the overall ability of CaLiV to compensate for the propagating S2V calibration error over time/distance. After performing the calibration with CaLiV, the validation target seen on the right side of Fig. 7 is well aligned, indicating a successful S2V calibration. To achieve these results, a number of criteria must be met in the real experiments. To account for the motion distortion of the LiDAR sensors and the distortion of the angular velocities during the curved trajectory, the vehicle motion must be stationary and at a constant velocity of about 1.4 $\frac{m}{s}$. At this velocity, a 20 Hz LiDAR sensor, and with the object covering only a small fraction of the LiDAR FoV, the effect of motion distortion can be mitigated for our experiments. Furthermore, real-world LiDAR sensors do not have homogeneous point resolution across the vertical and horizontal FoV, resulting in sparse point clouds at the edges of the FoV. In our experiments, this required an additional voxel and outlier filter to account for the differences in cardinality and improve the registration results during calibration.
@@ VI CONCLUSION & FUTURE WORK
The goal of this work was to provide a target-based framework for S2S and S2V calibration of multi-LiDAR systems with arbitrary FoV overlap. In addition, we aimed to develop an algorithm that is independent of environmental features and does not rely on external sensors. CaLiV demonstrates high accuracy and robustness for both S2S and S2V calibration. CaLiV outperforms the current state of the art in both non-overlapping sensor calibration and S2V calibration. Our current method requires a curved trajectory to optimize S2S and S2V calibration. For the S2S errors it is also possible to move the target instead of the vehicle and it remains to be investigated whether this would yield equally accurate results. Furthermore, the possibility of using a more complex trajectory could be explored to also solve the S2V translation errors while maintaining the accuracy.